A father is twice as old as his son. Ten years ago he was three times as old as his son.
a) If the son is x years old, determine the age of the father in terms of x.
b) Find the ages of the father and son ten years ago in terms of x.
c) What are their present ages.

Answer by ankor@dixie-net.com(15638)

A father is twice as old as his son. Ten years ago he was three times as old as his son.
:
a) If the son is x years old, determine the age of the father in terms of x.
f = 2x
:
b) Find the ages of the father and son ten years ago in terms of x.
f - 10 = 3(x - 10)
f - 10 = 3x - 30
f = 3x - 30 + 10
f = 3x - 20; father in terms of x, 10 years ago
:
c) What are their present ages.
replace f with 2x in the above equation
2x = 3x - 20
20 = 3x - 2x
20 = x
son is 20, then father is 40